# Distillation versus Contrastive Learning: How to Train Your Rerankers

### v2

@@ Abstract
This paper empirically compares these strategies by training rerankers of different sizes and architectures using both methods on the same data, with a strong contrastive learning model acting as the distillation teacher. Our results show that knowledge distillation generally yields better in-domain and out-of-domain ranking performance than contrastive learning when distilling from a larger teacher model. This finding is consistent across student model sizes and architectures. However, distilling from a teacher of the same capacity does not provide the same advantage, particularly for out-of-domain tasks. These findings offer practical guidance for choosing a training strategy based on available teacher models. We recommend using knowledge distillation to train smaller rerankers if a larger, more powerful teacher is accessible; in its absence, contrastive learning remains a robust baseline.
@@ 1 Introduction
Modern information retrieval (IR) systems often rely on a two-stage process: an initial retriever quickly finds candidate texts, and a more powerful reranker re-orders them by relevance to improve the final ranking quality (Schütze et al., 2008; Ma et al., 2023; Asai et al., 2024; Singh et al., 2025, inter alia). This reranking stage, which involves scoring and sorting texts by their relevance to a given query, commonly employs powerful cross-encoders Yates et al. (2021); Zhuang et al. (2024b).
@@
Effective training is key to building a high-performing cross-encoder reranker. Two main strategies have emerged for this purpose. The first, contrastive learning (CL), trains the model directly using ground-truth relevance labels, learning to distinguish positive (relevant) examples from negative (irrelevant) ones Oord et al. (2018); Gao et al. (2021); Ma et al. (2023). The second, knowledge distillation (KD), instead involves training a smaller “student” model to replicate the outputs of a larger, more capable “teacher” model Buciluǎ et al. (2006); Hinton et al. (2015); Hofstätter et al. (2020); Schlatt et al. (2025). This method is often employed to create efficient models that can approximate the performance of larger ones.
@@
Both strategies are widely used to train rerankers. But for practical deployments, is one preferable to the other? Prior work has explored related, but distinct, questions. For instance, Hofstätter et al. (2020) explored distilling from rerankers to retrievers; while Baldelli et al. (2024) and Schlatt et al. (2025) focused on training rerankers via distilling from large language models (LLMs) like GPT-4 Achiam et al. (2023). However, there is no direct comparison of these two strategies.
@@
In this paper, we address this gap. We present an empirical comparison of contrastive learning and knowledge distillation as training strategies for cross-encoder rerankers. To this end, we conduct a suite of controlled experiments. We train cross-encoder models of various sizes (0.5B, 1.5B, 3B, 7B) and different architectures (Transformer and Recurrent) using both strategies. For knowledge distillation, we employ a performant 7B Qwen2.5 model Yang et al. (2024) trained with contrastive learning as the teacher. We focus on modern decoder-only models, as recent works have reported their effectiveness over smaller encoder-based models like BERT (Ma et al., 2023; Muennighoff et al., 2024a; Zhang et al., 2025, inter alia). We train all models on the same dataset and evaluate their ranking performance on standard in-domain and out-of-domain benchmarks.
@@
Our findings consistently show that training with knowledge distillation yields better ranking performance than direct contrastive learning when the student model is smaller than the teacher. For instance, distilling from our 7B teacher significantly improved the performance of 0.5B, 1.5B, and 3B student models on both in-domain and out-of-domain benchmarks compared to training them with contrastive learning. This advantage holds across both Transformer and Recurrent model architectures. However, this benefit diminishes when the student and teacher have the same capacity; distilling from a 7B teacher to a 7B student yielded no significant improvements and, in some cases, harmed out-of-domain generalization Gou et al. (2021); Gholami and Omar (2023). To further test the robustness of our conclusions, we conduct an additional experiment where models are trained on a diverse, multi-domain dataset RLHN (Thakur et al., 2025) instead of only on hard negatives mined from a single source. In this setting, knowledge distillation again proves to be the more effective strategy, confirming its advantage even when the reranker’s training data is decoupled from the first-stage retriever. Together, these findings position knowledge distillation as a powerful and robust strategy for training smaller rerankers, provided a strong, larger teacher is available. In its absence, we find contrastive learning remains a strong baseline.
@@ 2 Methodology
Training strategy is critical for cross-encoder reranker’s ranking performance. Two primary strategies have been extensively studied in the literature: (1) direct optimization on ground-truth labels via contrastive learning Gao et al. (2021); Yates et al. (2021); Ma et al. (2023), and (2) knowledge transfer from a larger model via knowledge distillation (Hofstätter et al., 2020; Schlatt et al., 2025, inter alia).
@@
In the rest of this section, we first formally define the text reranking problem and our notation (§˜2.1). We then provide detailed technical descriptions of the training process using contrastive learning (§˜2.2) and knowledge distillation (§˜2.3).
@@ 2.1 The Text Reranking Problem
Modern IR systems often employ a two-stage retrieval-and-rerank pipeline (Schütze et al., 2008; Zhang et al., 2021; Asai et al., 2024, inter alia). An efficient first-stage retriever initially fetches a broad set of candidate texts. Subsequently, a more powerful reranker refines this initial list to optimize ranking metrics. Reranking is the task of ordering texts (e.g., passages or documents) by their relevance to a given query.
Let $q$ be an input query, and $d$ be a text from a corpus $\mathcal{D}$. We define a reranking model $f_{\theta}(q,d)$, parameterized by $\theta$, which computes a scalar relevance score. This model is typically a cross-encoder: the query $q$ and text $d$ are concatenated and fed into a transformer-based language model, whose output is passed through a linear layer to produce the final score (Yates et al., 2021; Nogueira et al., 2019; Boytsov et al., 2022; Ma et al., 2023; Xu, 2024; Xu et al., 2025b, inter alia).
@@ 2.2 Training with Contrastive Learning
Contrastive learning is one strategy for learning representations Oord et al. (2018); Weng (2021). Its application to reranking builds on the principle of the InfoNCE loss Oord et al. (2018), which is derived from Noise-Contrastive Estimation Gutmann and Hyvärinen (2010).
The general goal is to learn a model that distinguishes a “positive” data sample from a set of “negative” (or noise) samples, given a certain context. Denote a context vector $\mathbf{c}$, and consider a set of $N$ samples $X=\{\mathbf{x}_{i}\}_{i=1}^{N}$, where one sample $\mathbf{x}_{\text{pos}}$ is a positive sample drawn from the conditional distribution $p(\mathbf{x}|\mathbf{c})$, and the $N-1$ negative samples are drawn from a proposal distribution $p(\mathbf{x})$. Using Bayes’ rule, the probability that a sample $\mathbf{x}_{i}$ is the positive one is:
@@
$-\frac{1}{|\mathcal{S}|}\sum\limits_{(q_{i},d_{i}^{+})\in\mathcal{S}}\log\frac{\exp f_{\theta}(q_{i},d_{i}^{+})}{\exp f_{\theta}(q_{i},d_{i}^{+})+\sum\limits_{j\in\mathcal{D}_{i}^{-}}\exp f_{\theta}(q_{i},d_{i}^{-})}$
The total loss is averaged over all training instances. Following common practice Gao et al. (2021); Xu (2024), the negative documents are often "hard negatives" — documents that the first-stage retriever ranked highly but are not labeled as relevant. In practice, training instances are grouped into minibatches, and the parameters $\theta$ are optimized jointly.
@@ 2.3 Training with Knowledge Distillation
Knowledge distillation (KD) is a technique for training a smaller, efficient "student" model by transferring knowledge from a larger, more capable "teacher" model Hinton et al. (2015); Gou et al. (2021). In IR, KD is used to create fast rerankers that approximate the performance of slower, larger models, which is critical for production systems (Hofstätter et al., 2020; Santhanam et al., 2022, inter alia).
@@
where $T$ is the temperature hyperparameter. A higher temperature creates a softer probability distribution, which can help in transferring more nuanced information from the teacher. In practice, $T$ is often set to 1 Hinton et al. (2015).
@@ 3 Experimental Setup
We construct our training set based on the well established MS MARCO passage retrieval dataset (Bajaj et al., 2016). The official training set contains 532k training pairs, and the corpus contains 8.8M passages. For first stage retriever, we reproduce the RepLlama experiment Ma et al. (2023), but replace the Llama-2-7B backbone model Touvron et al. (2023) with stronger Qwen2.5-7B model Yang et al. (2024). The retriever — named RepQwen — is trained on Tevatron/MSMARCO-passage-aug [^1] https://huggingface.co/datasets/Tevatron/msmarco-passage-aug trainset, which consists of 486k $(q_{i},d_{i}^{+})$ pairs with hard negatives mined from BM25 and CoCondenser Gao and Callan (2022).
To construct the training set for the rerankers, we follow the same strategy as Ma et al. (2023) to mine hard negatives from the first stage retriever. Specifically, for each $(q_{i},d_{i}^{+})$ pair, we randomly sample $k$ negatives from top-200 passages retrieved by the retriever, excluding $d_{i}^{+}$. As shown in the experimental results (§˜4), the models trained using this constructed dataset achieves on par or improved performance over RankLlama2 Ma et al. (2023), suggesting the correctness of the pipeline. We did not further investigate hard negative mining strategies like prior works Yu et al. (2024); Thakur et al. (2025); Lee et al. (2025) as they are orthogonal to the goal of this paper.
@@
We evaluate the reranking performance following prior works’ practices (Ma et al., 2023; Xu et al., 2025c, inter alia). For in-domain evaluation, we use MS MARCO passage dev set, which includes 6,980 queries. We also include TREC DL19 and DL20 Craswell et al. (2020, 2021) consisting of 43 and 54 queries respectively. For out-of-domain evaluation, we adopt BEIR benchmark Thakur et al. (2021). We evaluate performances on 13 subsets that are publicly available. Refer to Appx.˜A for details of datasets.
@@
Our main experiments are based on RepQwen, the retriever with a Qwen2.5-7B backbone following the RepLlama Ma et al. (2023) training strategy. We also report the performance of a retriever with a Llama-3.1-8B backbone.
@@
We also experiment with RecurrentGemma De et al. (2024), a recurrent language model based on the Griffin architecture instead of quadratic complexity Transformers Vaswani et al. (2017). We use the 2B variant and refer to the trained model as RankRGemma-2B. Recent works have explored recurrent language models’ efficacy for IR tasks, such as state space models like Mamba Gu and Dao (2023); Dao and Gu (2024) for retrieval Zhang et al. (2024) and reranking Xu (2024); Xu et al. (2025c). We follow this direction to examine recurrent language models’ efficacy under different training strategies.
@@
Our baselines are prior results under the same training setting. RepLlama and RankLlama Ma et al. (2023) are the closest baselines which use the same trainset for contrastive learning, using Llama-2 backbone Touvron et al. (2023). CSPLADE Xu et al. (2025a) is a learned sparse retrieval model with Llama-3-8B backbone Dubey et al. (2024), and achieves competitive performance compared to dense retrieval models. RankMamba-2 Xu et al. (2025c) trains Mamba-2-based rerankers for passage reranking, though their results are based on BGE retriever Xiao et al. (2023). We compare against the original numbers reported by the authors.
@@
Our implementation is based on packages including PyTorch, Huggingface Transformers and Tevatron-v2 Ma et al. (2025). For all our models, we train with LoRA Hu et al. (2021) to balance in-domain and out-of-domain performance and reduce overfitting. For scalable training, we use DeepSpeed stage 2 Aminabadi et al. (2022), activation checkpointing, FlashAttention-2 Dao (2024), mixed precision and gradient accumulation. Appx.˜B gives details about hyperparameters.

### v3

@@ Abstract
This paper empirically compares these strategies by training rerankers of different sizes (0.5B, 1.5B, 3B, 7B) and architectures (Transformer, Recurrent) using both methods on the same data, with a strong contrastive learning model acting as the distillation teacher. Our results show that knowledge distillation generally yields better in-domain and out-of-domain ranking performance than contrastive learning when distilling from a more performant teacher model. This finding is consistent across student model sizes and architectures. However, distilling from a teacher of the same capacity does not provide the same advantage, particularly for out-of-domain tasks. These findings offer practical guidance for choosing a training strategy based on available teacher models. We recommend using knowledge distillation to train smaller rerankers if a larger, more performant teacher is accessible; in its absence, contrastive learning remains a robust baseline. Our code implementation is made available to facilitate reproducbility. [^1] https://github.com/zhichaoxu-shufe/distillation-versus-contrastive-learning
@@ 1 Introduction
Modern information retrieval (IR) systems often rely on a two-stage process: an initial retriever quickly finds candidate texts, and a more powerful reranker re-orders them by relevance to improve the final ranking quality (schutze2008introduction; ma2023fine; asai2024openscholar; singh2025ai2scholarqa, inter alia). This reranking stage, which involves scoring and sorting texts by their relevance to a given query, commonly employs powerful cross-encoders yates2021pretrained; zhuang2024setwiseapproachforeffectiveandhighlyefficient.
@@
Effective training is key to building a high-performing cross-encoder pointwise reranker. Two main strategies have emerged for this purpose. The first, contrastive learning (CL), trains the model directly using ground-truth relevance labels, learning to distinguish positive (relevant) examples from negative (irrelevant) ones oord2018representation; gao2021rethink; ma2023fine. The second, knowledge distillation (KD), instead involves training a smaller “student” model to replicate the outputs of a larger, more capable “teacher” model buciluǎ2006modelcompression; hinton2015distillingknowledge; hofstatter2020marginmse; schlatt2025rankdistillm. This method is often employed to create efficient models that can approximate the performance of larger ones.
@@
Both strategies are widely used to train rerankers. But for practical deployments, is one preferable to the other? Prior work has explored related, but distinct, questions. For instance, hofstatter2020marginmse explored distilling from rerankers to retrievers; while baldelli2024twolaratwostepllmaugmenteddistillation and schlatt2025rankdistillm focused on training rerankers via distilling from large language models (LLMs) like GPT-4 achiam2023gpt4techinicalreport. A line of works empirically compared distillation versus contrastive learning in training learning-to-rank models and BERT-sized rankers qin2021improvingneuralrankingvialosslessknowledgedistillation; qin2023rdsuite; qin2021neuralrankers; zhuang2021ensembledistillation. However, there lacks a direct comparison of these two strategies in the context of modern rerankers based on casual language models ma2023fine; xu-etal-2025-statespacemodelsarestrongtextrerankers.
@@
In this paper, we directly bridge this gap by presenting an empirical comparison of contrastive learning and knowledge distillation as training strategies for cross-encoder rerankers. To this end, we conduct a suite of controlled experiments. We train cross-encoder models of various sizes (0.5B, 1.5B, 3B, 7B) and different architectures (Transformer and Recurrent) using both strategies. For knowledge distillation, we employ a performant 7B Qwen2.5 model yang2024qwen2.5technicalreport trained with contrastive learning as the teacher. We focus on modern decoder-only models, as recent works have reported their effectiveness over smaller encoder-based models like BERT (ma2023fine; muennighoff2024gritlm; zhang2025qwen3embedding, inter alia). We train all models on the same dataset and evaluate their ranking performance on standard in-domain and out-of-domain benchmarks.
@@
Our findings consistently show that training with knowledge distillation yields better ranking performance than direct contrastive learning when the student model is smaller than the teacher. For instance, distilling from our 7B teacher significantly improved the performance of 0.5B, 1.5B, and 3B student models on both in-domain and out-of-domain benchmarks compared to training them with contrastive learning. This advantage holds across both Transformer and Recurrent model architectures. However, this benefit diminishes when the student and teacher have the same capacity; distilling from a 7B teacher to a 7B student yielded no significant improvements and, in some cases, harmed out-of-domain generalization gou2021knowledgedistillationsurvey; Gholami2023canastudentllmperformaswellasteacher. To further test the robustness of our conclusions, we conduct an additional experiment where models are trained on a diverse, multi-domain dataset RLHN (thakur2025fixingdatathathurtsperformance) instead of only on hard negatives mined from a single source. In this setting, knowledge distillation again proves to be the more effective strategy, confirming its advantage even when the reranker’s training data is decoupled from the first-stage retriever. Together, these findings position knowledge distillation as a powerful and robust strategy for training smaller rerankers, provided a performant teacher is available. In its absence, we find contrastive learning remains a robust baseline.
2 Related Works
In this paper, we focus exclusively on the architecture of cross-encoder pointwise reranker nogueira2019passage; nogueira2020document; zhuang2024setwiseapproachforeffectiveandhighlyefficient. The reranker model processes a query and a document simultaneously, allowing the self-attention mechanism to explicitly model the interactions between their tokens throughout the layers. nogueira2019passage first demonstrated the efficacy of this approach by fine-tuning BERT as a text pair classifier. The fine-tuning of the BERT-type model as cross-encoder reranker has later been extended to encoder-decoder architectures, such as T5 raffel2020exploring; nogueira2020document, and decoder-only models like Llama touvron2023llama; ma2023fine.
The training objective for a cross-encoder is to accurately discriminate between relevant and non-relevant documents for a given query. In practice, the non-relevant documents are usually hard negatives mined from first-stage retrievers (nogueira2019document; gao2021rethink; boytsov2022understanding, inter alia), and optionally combined with synthetic data curated from a complex and scalable data pipeline zhang2025qwen3embedding; wang2025jinarerankerv3. In literature, two primary training strategies are widely used to train cross-encoder pointwise reranker. The first is contrastive learning, which trains the model directly on ground-truth labels to maximize the distinction between positive and negative examples gao2021rethink; zhuang2023rankt5. The second strategy is knowledge distillation, which involves training a student reranker to mimic the behavior and performance of a larger, more capable teacher model hofstatter2020marginmse; baldelli2024twolaratwostepllmaugmenteddistillation; schlatt2025rankdistillm. Knowledge distillation (KD, buciluǎ2006modelcompression; hinton2015distillingknowledge) facilitates the transfer of knowledge from a large, complex “teacher” model to a smaller, more efficient “student” model. The goal is to enable the student to mimic the teacher’s output at the logit level, thereby inheriting its predictive capabilities at a reduced computational cost. Within the context of training neural IR models, knowledge distillation has been effectively applied to enhance both bi-encoder retriever (hofstatter2020marginmse; formal2021spladev2sparselexicalandexpansionmodel, inter alia) and cross-encoder reranker baldelli2024twolaratwostepllmaugmenteddistillation; schlatt2025rankdistillm. [^2] Note that, we use a narrower definition of knowledge distillation that focuses on aligning the student’s output with the teacher’s at the logit level hinton2015distillingknowledge. Sequence-level distillation kim2016sequencelevel, and training with synthetic data annotated by LLMs zhang2025qwen3embedding; schlatt2025rankdistillm fall into the broader scope of knowledge distillation. In this work, we aim to compare these two training strategies in a controlled experimental setting to evaluate their strengths and weaknesses.
@@ 3 Methodology
Training strategy is critical for cross-encoder reranker’s ranking performance. Two primary strategies have been extensively studied in the literature: (1) direct optimization on ground-truth labels via contrastive learning gao2021rethink; yates2021pretrained; ma2023fine, and (2) knowledge transfer from a larger model via knowledge distillation (hofstatter2020marginmse; schlatt2025rankdistillm, inter alia).
@@
In the rest of this section, we first formally define the text reranking problem and our notation (§˜3.1). We then provide detailed technical descriptions of the training process using contrastive learning (§˜3.2) and knowledge distillation (§˜3.3).
@@ 3.1 The Text Reranking Problem
Modern IR systems often employ a two-stage retrieval-and-rerank pipeline (schutze2008introduction; zhang-etal-2021-learning-rank; asai2024openscholar, inter alia). An efficient first-stage retriever initially fetches a broad set of candidate texts. Subsequently, a more powerful reranker refines this initial list to optimize ranking metrics. Reranking is the task of ordering texts (e.g., passages or documents) by their relevance to a given query.
Let $q$ be an input query, and $d$ be a text from a corpus $\mathcal{D}$. We define a reranking model $f_{\theta}(q,d)$, parameterized by $\theta$, which computes a scalar relevance score. This model is typically a cross-encoder: the query $q$ and text $d$ are concatenated and fed into a transformer-based language model, whose output is passed through a linear layer to produce the final score (yates2021pretrained; nogueira2019document; boytsov2022understanding; ma2023fine; xu2024rankmamba; xu2025surveyofmodelarchitectures, inter alia).
@@ 3.2 Training with Contrastive Learning
Contrastive learning is one strategy for learning representations oord2018representation; weng2021contrastive. Its application to reranking builds on the principle of the InfoNCE loss oord2018representation, which is derived from Noise-Contrastive Estimation gutmann10noisecontrastiveestimation.
The general goal is to learn a model that distinguishes a “positive” data sample from a set of “negative” (or noise) samples, given a certain context. Denote a context vector $\mathbf{c}$, and consider a set of $N$ samples $X=\{\mathbf{x}_{i}\}_{i=1}^{N}$, where one sample $\mathbf{x}_{\text{pos}}$ is a positive sample drawn from the conditional distribution $p(\mathbf{x}|\mathbf{c})$, and the N1N-1 negative samples are drawn from a proposal distribution $p(\mathbf{x})$. Using Bayes’ rule, the probability that a sample $\mathbf{x}_{i}$ is the positive one is:
@@
$\frac{1}{|\mathcal{S}|}\sum\limits_{(q_{i},d_{i}^{+})\in\mathcal{S}}\log\frac{\exp f_{\theta}(q_{i},d_{i}^{+})}{\exp f_{\theta}(q_{i},d_{i}^{+})+\sum\limits_{j\in\mathcal{D}_{i}^{-}}\exp f_{\theta}(q_{i},d_{i}^{-})}$
The total loss is averaged over all training instances. Following common practice gao2021rethink; xu2024rankmamba, the negative documents are often "hard negatives" — documents that the first-stage retriever ranked highly but are not labeled as relevant. In practice, training instances are grouped into minibatches, and the parameters $\theta$ are optimized jointly.
@@ 3.3 Training with Knowledge Distillation
Knowledge distillation (KD) is a technique for training a smaller, efficient "student" model by transferring knowledge from a larger, more capable "teacher" model hinton2015distillingknowledge; gou2021knowledgedistillationsurvey. In IR, KD is used to create fast rerankers that approximate the performance of slower, larger models, which is critical for production systems (hofstatter2020marginmse; santhanam-etal-2022-colbertv2, inter alia).
@@
where $T$ is the temperature hyperparameter. A higher temperature creates a softer probability distribution, which can help in transferring more nuanced information from the teacher. In practice, $T$ is often set to 1 hinton2015distillingknowledge.
@@ 4 Experimental Setup
We construct our training set based on the well established MS MARCO passage retrieval dataset (bajaj2016ms). The official training set contains 532k training pairs, and the corpus contains 8.8M passages. For first stage retriever, we reproduce the RepLlama experiment ma2023fine, but replace the Llama-2-7B backbone model touvron2023llama with stronger Qwen2.5-7B model yang2024qwen2.5technicalreport. The retriever — named RepQwen — is trained on Tevatron/MSMARCO-passage-aug [^3] https://huggingface.co/datasets/Tevatron/msmarco-passage-aug trainset, which consists of 486k $(q_{i},d_{i}^{+})$ pairs with hard negatives mined from BM25 and CoCondenser gao2022unsupervised.
To construct the training set for the rerankers, we follow the same strategy as ma2023fine to mine hard negatives from the first stage retriever. Specifically, for each $(q_{i},d_{i}^{+})$ pair, we randomly sample $k$ negatives from top-200 passages retrieved by the retriever, excluding $d_{i}^{+}$. As shown in the experimental results (§˜5), the models trained using this constructed dataset achieves on par or improved performance over RankLlama2 ma2023fine, suggesting the correctness of the pipeline. We did not further investigate hard negative mining strategies like prior works yu2024arcticembed2.0; thakur2025fixingdatathathurtsperformance; lee2025nvembed as they are orthogonal to the goal of this paper.
@@
We evaluate the reranking performance following prior works’ practices (ma2023fine; xu-etal-2025-statespacemodelsarestrongtextrerankers, inter alia). For in-domain evaluation, we use MS MARCO passage dev set, which includes 6,980 queries. We also include TREC DL19 and DL20 craswell2020overview; craswell2021overview consisting of 43 and 54 queries respectively. For out-of-domain evaluation, we adopt BEIR benchmark thakurbeir. We evaluate performances on 13 subsets that are publicly available. Refer to Appendix˜A for details of datasets.
@@
Our main experiments are based on RepQwen, the retriever with a Qwen2.5-7B backbone following the RepLlama ma2023fine training strategy. We also report the performance of a retriever with a Llama-3.1-8B backbone.
@@
We also experiment with RecurrentGemma de2024griffin, a recurrent language model based on the Griffin architecture instead of quadratic complexity Transformers vaswani2017attention. We use the 2B variant and refer to the trained model as RankRGemma-2B. Recent works have explored recurrent language models’ efficacy for IR tasks, such as state space models like Mamba gu2023mamba; dao2024transformers for retrieval zhang2024mamba and reranking xu2024rankmamba; xu-etal-2025-statespacemodelsarestrongtextrerankers. We follow this direction to examine recurrent language models’ efficacy under different training strategies.
@@
Our baselines are prior results under the same training setting. RepLlama and RankLlama ma2023fine are the closest baselines which use the same trainset for contrastive learning, using Llama-2 backbone touvron2023llama. CSPLADE xu2025csplade is a learned sparse retrieval model with Llama-3-8B backbone dubey2024llama, and achieves competitive performance compared to dense retrieval models. RankMamba-2 xu-etal-2025-statespacemodelsarestrongtextrerankers trains Mamba-2-based rerankers for passage reranking, though their results are based on BGE retriever bge_embedding. We compare against the original numbers reported by the authors.
@@
Our implementation is based on packages including PyTorch, Huggingface Transformers and Tevatron-v2 ma2025tevatron. For all our models, we train with LoRA hu2021lora to balance in-domain and out-of-domain performance and reduce overfitting. For scalable training, we use DeepSpeed stage 2 aminabadi2022deepspeed, activation checkpointing, FlashAttention-2 dao2024flashattention, mixed precision and gradient accumulation. Appendix˜B gives details about hyperparameters.